q=100+p-0.01p2+2n-0.03n2

determine the marginal rate of technical substitution.
1
Expert's answer
2011-12-06T11:01:07-0500
Marginal Rate of Technical Substitution (or Technical Rate of Substitution) is the amount by which the quantity of one input (n)& has to be reduced when one extra unit of another input (n2) is used, so that the total output (q) remains constant:

Dn/Dn2 = MP(n2)/MP(n) = -0.03/2= -0.015 – MRTS for the input n for n2